An X-ray photon has a momentum of 4.2 x 10-23 kg m/s. Determine its energy in keV.

Your answer should be a number with one decimal place.
Answer #1

P = 4.2×10^-23kg-m/s

Since P = h/lemda

=) Lemda = h/p = 6.63×10^-34/4.2×10^-23 = 1.578×10^-11m

=) Wavelength lemda = 1.587×10^-11 m = 0.01587nm

Now energy E = hc/lemda = 1240eV.nm/0.01587nm

=) E = 78.2 keV